OpenStudy (amtran_bus):

Is there a way to get wolfram to show area of region? Trying to figure out how to use it.

OpenStudy (phi):

Does not look promising... is this an integral problem?

OpenStudy (amtran_bus):

Well, yes and no. Here is all I know.

OpenStudy (amtran_bus):

Along with choices Area = 14.79 Area = 14.77 Area = 14.78 Area = 14.98

OpenStudy (phi):

then do \[ \int_{-4}^\infty e^{-\frac{x}{2}} \ dx\] if you know how do you know calculus?

OpenStudy (amtran_bus):

For real? Is that it? I was so overcomplicating this. I can do that.

OpenStudy (amtran_bus):

2e^2

OpenStudy (phi):

let u = -x/2 du = -dx/2 so dx = -2 du u goes from -(-4)/2 = 2 to u = -infinity/2 = -infinity thus \[ - 2 \int_2^{-\infty} e^u \ du \] or if we negate the integral we can swap the order of the integrand \[ 2 \int_{-\infty}^2 e^u \ du = 2 e^u \bigg|_{-\infty}^2= 2e^2 - 0= 2 e^2\]

OpenStudy (amtran_bus):

Phi, you are a lifesaver. Thanks a million.